OpenStudy (anonymous):

Find the product. (0.5n^5)^2* (10n^7)^3 Answer is something like... _ n to the power of _ The amount of underscores have nothing to do with the number of digits. Just means there's a blank.

OpenStudy (bahrom7893):

0.25n^10 * 1000*n^21 = 0.25*1000 * n^(10+21)

OpenStudy (anonymous):

Uhm, I don't think that's right. Doesn't seem so.

OpenStudy (bahrom7893):

0.5*n^5 * 0.5*n^5 * 10n^7 * 10n^7 * 10n^7

OpenStudy (anonymous):

Hmm... Okay. Thanks.

OpenStudy (bahrom7893):

SO now just multiply it out using the rules of exponents: 0.25*n^(5+5) * 1000*n^(7+7+7) = 0.25 * n^10 * 1000 * n^21

OpenStudy (bahrom7893):

= 0.25*1000 * n^(10+21) = 250 * n^31

OpenStudy (anonymous):

Thank you :)